OpenStudy (anonymous):

Which of the following represents eighth root of x to the fifth power in exponential form? 5x8 8x5 x to the five eighths power x to the eight fifths power

OpenStudy (anonymous):

\[\sqrt[8]{x^5} \] A: \[5x^8 \] B: \[8x^5 \] C: \[x ^{5/8} \] D: \[x ^{8/5}\]

OpenStudy (anonymous):

I have no clue

OpenStudy (anonymous):

explain?

OpenStudy (phi):

when you see "root" as in eighth root think "exponent is 1/ 8" so eight root of x can be written as \[ \sqrt[8]{x} = x^\frac{1}{8} \] to raise it to the 5th power i.e. exponent = 5 write \[ \left(x^\frac{1}{8} \right)^5 \] use the rule \[ \left(a^b\right)^c = a^{bc} \] to simplify to \[ x^\frac{5}{8} \]

OpenStudy (phi):

so 3 things to know: "nth root" means EXPONENT 1/n raise to the power m means EXPONENT m and the rule \[ \left(a^\frac{1}{n}\right)^m = a^{\frac{1}{n}\cdot m}= a^\frac{m}{n} \]

OpenStudy (anonymous):

Thanks